# Assimilative Causal Inference

### v1

@@ 3.1 Online Smoothing for Dynamic Influence Tracking
Recall from (2.4) that the relative entropy $\mathcal{P}(p_{t}^{\text{s}},p_{t}^{\text{f}})$ between the filter and smoother estimates quantifies the causal influence from $\mathbf{y}$ to $\mathbf{x}$. Here, $p_{t}^{\text{s}}=p(\mathbf{y}(t)|\mathbf{x}(s\leq T))$ incorporates the complete future information of $\mathbf{x}$. Due to the finite effective influence length, the complete smoother solution $p_{t}^{\text{s}}$ achieves a substantial uncertainty reduction compared to the lagged posterior distributions $p(\mathbf{y}(t)|\mathbf{x}(s\leq T^{\prime}))$, which only utilize future information up to $t\leq T^{\prime}\leq T$, for a short time period after $t$. The divergence between these distributions vanishes as $T^{\prime}\rightarrow T$. We therefore define the causal influence radius (CIR) through the relative entropy criterion: For a predetermined threshold $\varepsilon>0$, the maximum lag $T^{\prime}_{\varepsilon}$ satisfying
$\mathcal{P}\left(p(\mathbf{y}(t)|\mathbf{x}(s\leq T)),p(\mathbf{y}(t)|\mathbf{x}(s\leq T^{\prime}))\right)>\varepsilon,$ (3.1)
yields the innate CIR measure $\tau_{\varepsilon}(t)=T^{\prime}_{\varepsilon}-t$. This adaptive time variable $\tau_{\varepsilon}(t)$ defines the effective causal period $[t,t+\tau_{\varepsilon}(t)]$ during which $\mathbf{y}(t)$ exerts measurable influence on $\mathbf{x}$. A recently developed online smoother algorithm [1] is employed to compute the CIR. This online smoother computes the uncertainty reduction in the estimated state of $\mathbf{y}(t)$ when the future information of $\mathbf{x}$ is sequentially added. See the Supporting Information for more technical details.
While the CIR computed via (3.1) is intuitive, it remains inherently subjective due to its dependence on the threshold parameter $\epsilon$. In the following, an objective CIR formulation is developed, eliminating the need for a prescribed threshold. This threshold-independent measure is subsequently integrated into the ACI framework.
@@ 3.2 Objective CIR
For simplicity, assume that the relative entropy between the aforementioned complete and lagged smoother distributions decreases monotonically as more future information of $\mathbf{x}$ is incorporated. This commonly holds in practice, though the theory of the objective CIR (developed rigorously in the Supporting Information) doesn’t require this condition. At a given time instant $t$, let $M=\mathcal{P}(p_{t}^{\text{s}},p_{t}^{\text{f}})$ denote the maximum value of the relative entropy between the complete and lagged smoother distributions over $t\leq T^{\prime}\leq T$, which is the difference between the smoother estimate and the filter estimate (where the latter can be viewed as a degenerate smoother using no future information). Given a threshold $\varepsilon\in[0,M]$, (3.1) yields a subjective CIR as $\tau_{\varepsilon}(t)=T^{\prime}_{\varepsilon}-t$. The objective CIR is then naturally defined by
[…]
Dividing $M$ guarantees the unit of the objective CIR is ‘time’. See Panel (d) of Figure 2.1 for a schematic illustration of the objective CIR and its relationship with the subjective ones. Note that numerically evaluating the integral in (3.2) requires repeated smoother computations, leading to a computational complexity that scales quadratically with the number of discretization points. An efficient computational method of (3.2) is included in the Supporting Information.
@@ 4 Conditional ACI in the Presence of Additional Non-Target Variables
In the more general context of a complex system, state variables can be grouped into three disjoint subsets: $\mathbf{x}_{\text{A}}$, $\mathbf{x}_{\text{B}}$, and $\mathbf{y}$. The goal is to determine whether $\mathbf{y}$ is a cause of $\mathbf{x}_{\text{A}}$ in the presence of the remaining variables $\mathbf{x}_{\text{B}}$, which are called the non-target variables, including the cases of confounders or mediators. Since $\mathbf{x}_{\text{B}}$ interacts with both $\mathbf{x}_{\text{A}}$ and $\mathbf{y}$, appropriately accounting for the role of $\mathbf{x}_{\text{B}}$ while inferring the causal link from $\mathbf{y}$ to $\mathbf{x}_{\text{A}}$ becomes essential. Hereafter, let us assume that a realization for each of $\mathbf{x}_{\text{A}}$ and $\mathbf{x}_{\text{B}}$, namely, $\mathbf{x}_{\text{A}}(s\leq T)$ and $\mathbf{x}_{\text{B}}(s\leq T)$ is available. Therefore, the candidate cause $\mathbf{y}$ is the only variable that does not require to have observations in this framework. Following the argument in Section 2, the causal relationship from $\mathbf{y}(t)$ to $(\mathbf{x}_{\text{A}},\mathbf{x}_{\text{B}})$ is identified based on the uncertainty reduction in the smoother distribution $p_{t}^{\text{s}}$ related to the filter one $p_{t}^{\text{f}}$, where
$\begin{split}p_{t}^{\text{f}}&=p(\mathbf{y}(t)|\mathbf{x}_{\text{A}}(s\leq t),\mathbf{x}_{\text{B}}(s\leq t)),\\ p_{t}^{\text{s}}&=p(\mathbf{y}(t)|\mathbf{x}_{\text{A}}(s\leq T),\mathbf{x}_{\text{B}}(s\leq T)).\end{split}$ (4.1)
@@
The main focus here is to appropriately determine $p_{t}^{\text{f}}$ and $p_{t}^{\text{s}}$ in the presence of $\mathbf{x}_{\text{B}}$.
@@
Recall from Section 2 that ACI employs Bayesian data assimilation to estimate the state of $\mathbf{y}$. ACI exploits observations of $\mathbf{x}$ to reduce the uncertainty in $\mathbf{y}$, where the influence of $\mathbf{x}$ on $\mathbf{y}$ depends on its own uncertainty: smaller observational uncertainty leads to a stronger impact. When extending this framework to include $\mathbf{x}_{\text{B}}$, the observed $\mathbf{x}_{\text{B}}$ time series is still used in the forecast step to compute uncertainties in $\mathbf{x}_{\text{A}}$ and $\mathbf{y}$. However, during the analysis step, the uncertainty in $\mathbf{x}_{\text{B}}$’s dynamics is deliberately excluded by assigning infinite uncertainty to its likelihood before posterior computation. Inflating the uncertainty in $\mathbf{x}_{\text{B}}$ is straightforward using Bayesian inference for state estimation. This ensures that $\mathbf{x}_{\text{B}}$ does not affect the uncertainty reduction in the state estimation of $\mathbf{y}$ when treated as an observational process. Consequently, any uncertainty reduction in $\mathbf{y}$ results solely from $\mathbf{x}_{\text{A}}$ under this proposed pipeline. See Panel (c) of Figure 2.1 for an illustration. For dynamical systems with explicit governing equations, a shortcut to exclude the influence in the uncertainty of $\mathbf{x}_{\text{B}}$ is to treat $\mathbf{x}_{\text{B}}$ as a prescribed forcing term in the reduced system defined by $(\mathbf{x}_{\text{A}},\mathbf{y})$ during Bayesian inference, with values defined by its observed time series system. Nevertheless, this framework does not necessarily require modifications to the underlying dynamical system, making it compatible with any given model, potentially including operational models in atmospheric and ocean science, as the model structure and integrity remains intact.
@@
It is important to emphasize that while treating $\mathbf{x}_{\text{B}}$ as a conditioning variable may appear analogous to the role of non-target variables in transfer entropy and other traditional causal inference methods, the way of handling the uncertainty in ACI is fundamentally different. Conventional methods treat $\mathbf{x}_{\text{B}}$ as a fixed component. In contrast, ACI explicitly accounts for $\mathbf{x}_{\text{B}}$’s dynamical influence while systematically prohibiting its uncertainty from affecting state estimation. This unique treatment allows the inferred causal relationships to reflect true dynamical interactions by isolating the contribution of uncertainty reduction from the non-target variables. As a final remark, if Bayesian data assimilation is used to compute $p(\mathbf{y}(t),\mathbf{x}_{\text{B}}(t)|\mathbf{x}(s\leq T))$ and then marginalize over $\mathbf{x}_{\text{B}}(t)$ to obtain $p(\mathbf{y}(t)|\mathbf{x}(s\leq T))$, this manipulation may lead to incorrect conclusions. This is because the correlation between $\mathbf{y}(t)$ and $\mathbf{x}_{\text{B}}(t)$ unavoidably alters the uncertainty in $\mathbf{y}(t)$ during Bayesian data assimilation, potentially introducing a spurious causal relationship between $\mathbf{y}$ and $\mathbf{x}_{\text{A}}$. This is true even if we condition over $\mathbf{x}_{\text{B}}(t)$. For instance, for the causal chain $\mathbf{x}_{\text{A}}\to\mathbf{x}_{\text{B}}\to\mathbf{y}$, such a method might falsely imply a direct causal link between $\mathbf{x}_{\text{A}}$ and $\mathbf{y}$.
@@
The details of conditional ACI implementation, along with validation of (conditional) ACI’s nil causality principles, are provided in the Supporting Information.
@@ 5.1 A Nonlinear Dyad Model with Extreme Events
$\frac{{\mathrm{d}}y}{{\mathrm{d}}t}$ $=-d_{y}y-\gamma x^{2}+f_{y}+\sigma_{y}\dot{W}_{y}.$ (5.1b)
This is a reduced-order conceptual model for atmospheric variability. It has been used to analyze the effects of various coarse-grained procedures on processes exhibiting intermittency, large-scale bifurcations, and microscale phase transitions. It is defined by an energy-conserving condition on its quadratic nonlinearities [25]. The following parameter values are used for this model:
@@
Figure 5.1 illustrates the ACI and CIRs from $y$ to $x$. Panel (a) shows that extreme events in $x$ occur intermittently. When the combined coefficient $-d_{x}+\gamma y$ in (5.1a) becomes positive, the dynamics of $x$ exhibit anti-damping, leading to amplitude growth and extreme events. Conversely, when the coefficient is negative, $x$ behaves as a damped system, with fluctuations primarily driven by random noise. Panel (b) displays the ACI, revealing that phases with significant ACI values largely coincide with the onset and peak phases of extreme events in $x$. This aligns with intuition, as $y$ acts as an anti-damping source and the primary driver of these events. Once $x$ peaks and begins to decay, the ACI value drops sharply, reflecting the strong negative feedback from $x$ to $y$, which then dampens $x$. The whisker plot in Panel (a) highlights the objective CIR, indicating a sustained influence from $y$ to $x$ during the triggering phase of extreme events but minimal influence during their demise. Notably, the long-range influence from the CIR reveals that extreme events develop gradually, with triggering conditions established well in advance. From a data assimilation perspective, when the observable signal $x$ starts to strengthen, it enhances the ability of the smoother to estimate $y$ as the signal in the future contains useful information beyond noise. As a result, a longer CIR is obtained. In contrast, once $x$ becomes sufficiently strong, its high signal-to-noise ratio ensures the filter captures all relevant triggering dynamics, leaving little room for the smoother to improve estimates. Hence, the outset of the ACI metric’s decline and shorter CIR. The Supporting Information contains the filter and smoother distributions and the subjective CIR with different threshold values $\epsilon$.
Figure 5.1: ACI values and CIRs for the nonlinear dyad model (5.1) from $y$ to $x$ at each time instant $t$. Panel (a): Time series of $x$ (magenta) and $y$ (blue), the objective CIR represented by a whisker plot emanating forward in time from each $y(t)$ value, and the anti-damping threshold line at $d_{x}/\gamma$. Panel (b): ACI from $y$ to $x$ as a function of time.

### v2

@@ 3.1 Online Smoothing for Dynamic Influence Tracking
Recall from (2.5) that the relative entropy $\mathcal{P}(p_{t}^{\text{s}},p_{t}^{\text{f}})$ between the filter and smoother estimates quantifies the causal influence from $\mathbf{y}(t)$ to $\mathbf{x}$. Here, $p_{t}^{\text{s}}=p(\mathbf{y}(t)|\mathbf{x}(s\leq T))$ incorporates the complete future information of $\mathbf{x}$. Due to the finite effective influence length, the complete smoother solution $p_{t}^{\text{s}}$ achieves a substantial uncertainty reduction compared to the lagged posterior distributions $p(\mathbf{y}(t)|\mathbf{x}(s\leq T^{\prime}))$, which only utilize future information up to $t\leq T^{\prime}\leq T$, for a short time period after $t$. The divergence between these distributions vanishes as $T^{\prime}\rightarrow T$. We therefore define the causal influence radius or CIR through the relative entropy criterion: For a predetermined threshold $\varepsilon\geq 0$, the maximum lag $T^{\prime}_{\varepsilon}$ satisfying
$\mathcal{P}\big(p(\mathbf{y}(t)|\mathbf{x}(s\leq T)),p(\mathbf{y}(t)|\mathbf{x}(s\leq T^{\prime}))\big)>\varepsilon,$ (3.1)
yields the innate CIR measure $\uptau_{\varepsilon}(t)=T^{\prime}_{\varepsilon}-t$. This adaptive time variable $\uptau_{\varepsilon}(t)$ defines the effective causal period $t,t+\uptau_{\varepsilon}(t)]$ during which $\mathbf{y}(t)$ exerts measurable influence on $\mathbf{x}$. A recently developed online smoother algorithm [andreou2024adaptive] is employed to compute the CIR. This online smoother computes the uncertainty reduction in the estimated state of $\mathbf{y}(t)$ when the future information of $\mathbf{x}$ is sequentially added. See the Supplementary Information for more technical details.
While the CIR computed via (3.1) is intuitive, it remains inherently subjective due to its dependence on the threshold parameter $\varepsilon$. In the following, an objective CIR formulation is developed, eliminating the need for a prescribed threshold. This threshold-independent measure is subsequently integrated into the ACI framework.
@@ 3.2 Objective CIR
For simplicity, assume that the relative entropy between the aforementioned complete and lagged smoother distributions decreases monotonically as more future information of $\mathbf{x}$ is incorporated. This commonly holds in practice, though the theory of the objective CIR (developed rigorously in the Supplementary Information) doesn’t require this condition. At a given time instant $t$, let $M=\mathcal{P}(p_{t}^{\text{s}},p_{t}^{\text{f}})$ denote the maximum value of the relative entropy between the complete and lagged smoother distributions over $t\leq T^{\prime}\leq T$, which is the difference between the smoother estimate and the filter estimate (where the latter can be viewed as a degenerate smoother using no future information). Given a threshold $\varepsilon\in[0,M]$, (3.1) yields a subjective CIR as $\uptau_{\varepsilon}(t)=T^{\prime}_{\varepsilon}-t$. The objective CIR is then naturally defined by
$\uptau(t):=\frac{1}{M}\int_{0}^{M}\uptau_{\varepsilon}(t)d\varepsilon.$ (3.2)
Dividing $M$ guarantees the unit of the objective CIR is “time”. See Panel (d) of Figure 1 for a schematic illustration of the objective CIR and its relationship with the subjective ones. Note that numerically evaluating the integral in (3.2) requires repeated smoother computations, leading to a computational complexity that scales quadratically with the number of discretization points. An efficient computational method of (3.2) is included in the Supplementary Information.
@@ 4 Conditional ACI in the Presence of Additional Non-Target Variables
In the more general context of a complex system, state variables can be grouped into three disjoint subsets: $\mathbf{x}_{\text{A}}$, $\mathbf{x}_{\text{B}}$, and $\mathbf{y}$, with $\mathbf{\mathbf{x}}=(\mathbf{x}_{\text{A}},\mathbf{x}_{\text{B}})^{\mathtt{T}}$ under the previous notation. The goal is to determine whether $\mathbf{y}$ is a cause of $\mathbf{x}_{\text{A}}$ in the presence of the remaining variables $\mathbf{x}_{\text{B}}$, which are called the non-target variables, including the cases of confounders or mediators. Since $\mathbf{x}_{\text{B}}$ interacts with both $\mathbf{x}_{\text{A}}$ and $\mathbf{y}$, appropriately accounting for the role of $\mathbf{x}_{\text{B}}$ while inferring the causal link from $\mathbf{y}$ to $\mathbf{x}_{\text{A}}$ becomes essential. Hereafter, let us assume that a realization for each of $\mathbf{x}_{\text{A}}$ and $\mathbf{x}_{\text{B}}$, namely $\mathbf{x}_{\text{A}}(s\leq T)$ and $\mathbf{x}_{\text{B}}(s\leq T)$, is available. Therefore, the candidate cause $\mathbf{y}$ is the only variable that does not require to have observations in this framework. Following the argument in Section 2, the causal relationship from $\mathbf{y}(t)$ to $(\mathbf{x}_{\text{A}},\mathbf{x}_{\text{B}})$ is identified based on the uncertainty reduction in the smoother distribution $p_{t}^{\text{s}}$ related to the filter one $p_{t}^{\text{f}}$, where
$\begin{split}p_{t}^{\text{f}}&=p(\mathbf{y}(t)|\mathbf{x}_{\text{A}}(s\leq t),\mathbf{x}_{\text{B}}(s\leq t)),\\ p_{t}^{\text{s}}&=p(\mathbf{y}(t)|\mathbf{x}_{\text{A}}(s\leq T),\mathbf{x}_{\text{B}}(s\leq T)).\end{split}$
@@
The main focus here is to appropriately determine and resolve the presence of $\mathbf{x}_{\text{B}}$ from $p_{t}^{\text{f}}$ and $p_{t}^{\text{s}}$ as to infer a conditional causal relationship from $\mathbf{y}(t)$ to $\mathbf{x}_{\text{A}}$ beyond the contributions of $\mathbf{x}_{\text{B}}$.
@@
Recall from Section 2 that ACI employs Bayesian data assimilation to estimate the state of $\mathbf{y}(t)$. ACI exploits observations of $\mathbf{x}$ to reduce the uncertainty in $\mathbf{y}(t)$, where the influence of $\mathbf{x}$ on the estimation of $\mathbf{y}(t)$ depends on its own uncertainty: smaller observational uncertainty leads to a stronger impact [rozovsky2012stochastic]. When extending this framework to $\mathbf{\mathbf{x}}=(\mathbf{x}_{\text{A}},\mathbf{x}_{\text{B}})^{\mathtt{T}}$ for conditional ACI, the observed time series of $\mathbf{x}_{\text{B}}$ is still used in the forecast step to compute the uncertainties in $\mathbf{x}_{\text{A}}$ and $\mathbf{y}$. However, during the analysis step, the uncertainty in $\mathbf{x}_{\text{B}}$’s dynamics is deliberately excluded by assigning infinite uncertainty to its marginal likelihood before posterior computation. This can be rigorously formulated through a limit operation on the analysis component of the Bayesian data assimilation update equations, which formally leads to the following filter- and smoother-based distributions:
$p^{\text{f}|\mathbf{x}_{\text{B}}}_{t}:=\lim_{\mathrm{Var}(\mathbf{x}_{\text{B}}(t))\to+\infty}p_{t}^{\text{f}},\qquad p^{\text{s}|\mathbf{x}_{\text{B}}}_{t}:=\lim_{\mathrm{Var}(\mathbf{x}_{\text{B}}(t))\to+\infty}p_{t}^{\text{s}}.$
Similar to (2.5), if the relative entropy of $p^{\text{s}|\mathbf{x}_{\text{B}}}_{t}$ from $p^{\text{f}|\mathbf{x}_{\text{B}}}_{t}$ is positive, $\mathcal{P}(p^{\text{s}|\mathbf{x}_{\text{B}}}_{t},p^{\text{f}|\mathbf{x}_{\text{B}}}_{t})>0$, then $\mathbf{y}$ is identified as the cause of $\mathbf{x}_{\text{A}}$ conditional to $\mathbf{x}_{\text{B}}$ at time $t$. This means future data of $\mathbf{x}_{\text{A}}$ incur additional information gain about $\mathbf{y}(t)$ related to its history, beyond the observational contributions of $\mathbf{x}_{\text{B}}$. By computing this metric at each $t$, a time-dependent conditional causal relationship is established.
Inflating the uncertainty in $\mathbf{x}_{\text{B}}$ is straightforward using Bayesian inference for state estimation. This approach ensures that $\mathbf{x}_{\text{B}}$ does not affect the uncertainty reduction in the state estimation of $\mathbf{y}(t)$ when treated as an observational process. Consequently, any uncertainty reduction in $\mathbf{y}(t)$ results solely from $\mathbf{x}_{\text{A}}$ under this proposed pipeline. See Panel (c) of Figure 1 for an illustration. For dynamical systems with explicit governing equations (see (2.1)), a shortcut to exclude the influence in the uncertainty of $\mathbf{x}_{\text{B}}$ is to treat $\mathbf{x}_{\text{B}}$ as a prescribed forcing term in the reduced system defined by $(\mathbf{x}_{\text{A}},\mathbf{y})$ during Bayesian inference, with values defined by its observed time series. Nevertheless, this framework does not necessarily require modifications to the underlying dynamical system, making it compatible with any given model, potentially including operational models in atmospheric and ocean science, as the model structure and integrity remains intact.
@@
It is important to emphasize that while treating $\mathbf{x}_{\text{B}}$ as a conditioning variable may appear analogous to the role of non-target variables in transfer entropy and other traditional causal inference methods, the way of handling the uncertainty in ACI is fundamentally different. Conventional methods treat $\mathbf{x}_{\text{B}}$ as a fixed component. In contrast, ACI explicitly accounts for $\mathbf{x}_{\text{B}}$’s dynamical influence while systematically prohibiting its uncertainty from affecting state estimation. This unique treatment allows the inferred causal relationships to reflect true dynamical interactions by isolating the contribution of uncertainty reduction from the non-target variables. As a final remark, if Bayesian data assimilation is used to compute $p(\mathbf{y}(t),\mathbf{x}_{\text{B}}(t)|\mathbf{x}_{\text{A}}(s\leq T))$ and then marginalize over $\mathbf{x}_{\text{B}}(t)$ to obtain $p(\mathbf{y}(t)|\mathbf{x}_{\text{A}}(s\leq T))$, this manipulation may lead to incorrect conclusions. This is because the correlation between $\mathbf{y}(t)$ and $\mathbf{x}_{\text{B}}(t)$ unavoidably alters the uncertainty in $\mathbf{y}(t)$ during Bayesian data assimilation, potentially introducing a spurious causal relationship between $\mathbf{y}$ and $\mathbf{x}_{\text{A}}$. This is true even if we condition over $\mathbf{x}_{\text{B}}(t)$ to obtain $p(\mathbf{y}(t)|\mathbf{x}_{\text{A}}(s\leq T),\mathbf{x}_{\text{B}}(t))$. For instance, for the causal chain $\mathbf{y}\to\mathbf{x}_{\text{B}}\to\mathbf{x}_{\text{A}}$, such a method might falsely imply a direct causal link from $\mathbf{y}(t)$ to $\mathbf{x}_{\text{A}}$.
@@
The details of the conditional ACI implementation, along with the verification of (conditional) nil causality principles under ACI, are provided in the Supplementary Information.
@@ 5 Applications to Nonlinear Systems with Intermittency and Extreme Events
We apply the ACI framework to identify causal relationships across the state space and evaluate the corresponding CIRs in complex nonlinear dynamical systems exhibiting intermittency and extreme events. The dynamical properties of these systems and the mechanisms that trigger crucial observed phenomena are well understood. This knowledge is exploited as guidance to validate the results when applying ACI. In addition, a perfect-model setup is used in the case studies carried out, meaning the system generates the synthetic observational data and the forecast model in ACI at the same. This facilitates the recovery of causal relationships in the absence of model error. Because the underlying system is chaotic and subject to stochastic forcing, a perfect model means that the model used in ACI shares the same governing equations and long-term statistical behavior as the system that generates the true signal. Additional results, including further numerical experiments, case studies, and one crucial application using sparse real-world observations, are provided in the Supplementary Information.
The operational outline of the ACI metric used to infer instantaneous causal relationships in the following experiments, including the details of how we determine the associated CIRs and account for non-target variables, is described in Sections 24. The developed framework is also summarized via the schematic diagram provided in Figure 1. Additional mathematical results and analyses, as well as computational implementation details, can be found in the Supplementary Information.
@@ 5.1 A Nonlinear Dyad Model with Extreme Events
$\frac{{\mathrm{d}}y}{{\mathrm{d}}t}$ $=-d_{y}y-\gamma x^{2}+f_{y}+\sigma_{y}\dot{W}_{y},$ (5.1b)
where $\dot{W}_{x}$ and $\dot{W}_{y}$ are independent standardized normal random variables. Within the proposed ACI framework, only $x$ is assumed to be observed over time. This is a reduced-order conceptual model for atmospheric variability. It has been used to analyze the effects of various coarse-grained procedures on processes exhibiting intermittency, large-scale bifurcations, and microscale phase transitions. It is defined by an energy-conserving condition on its quadratic nonlinearities [majda2012physics]. The following parameter values are used for this model:
@@
Figure 2 illustrates the ACI and CIRs from $y$ to $x$. Panel (a) shows that extreme events in $x$ occur intermittently. When the combined coefficient $d_{x}+\gamma y$ in (5.1a) becomes positive, the dynamics of $x$ exhibit anti-damping, leading to amplitude growth and extreme events. Conversely, when the coefficient is negative, $x$ behaves as a damped system, with fluctuations primarily driven by random noise. Panel (b) displays the ACI, revealing that phases with significant ACI values largely coincide with the onset and peak phases of extreme events in $x$. This aligns with intuition, as $y$ acts as an anti-damping source and the primary driver of these events. Once $x$ peaks and begins to decay, the ACI value drops sharply, reflecting the strong negative feedback from $x$ to $y$, which then dampens $x$. The whisker plot in Panel (a) highlights the objective CIR, indicating a sustained influence from $y$ to $x$ during the triggering phase of extreme events but minimal influence during their demise. Notably, the long-range influence from the CIR reveals that extreme events develop gradually, with triggering conditions established well in advance. From a data assimilation perspective, when the observable signal $x$ starts to strengthen, it enhances the ability of the smoother to estimate $y$ as the signal in the future contains useful information beyond noise. As a result, a longer CIR is obtained. In contrast, once $x$ becomes sufficiently strong, its high signal-to-noise ratio ensures the filter captures all relevant triggering dynamics, leaving little room for the smoother to improve estimates. Hence, the outset of the ACI metric’s decline and shorter CIR. The Supplementary Information contains the filter and smoother distributions and the subjective CIR with different threshold values $\varepsilon$.
Figure 2: ACI values and CIRs from $y$ to $x$ at each time instant $t$ for the nonlinear dyad model (5.1). Panel (a): Time series of $x$ (magenta) and $y$ (blue), the objective CIR represented by a whisker plot emanating forward in time from each $y(t)$ value, and the anti-damping threshold line at $d_{x}/\gamma$. Panel (b): ACI from $y$ to $x$ as a function of time.